40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Apr 29 22:53:06 2018, gautam, Update, General, DARM actuation estimate darmEst.pdf
    Reply  Tue May 1 19:37:50 2018, gautam, Update, General, DARM actuation estimate darmEst.pdfdarmEst_time.pdf
       Reply  Mon May 7 16:23:06 2018, gautam, Update, General, DARM actuation estimate darmProj.pdf37.pngMCalign_20180507.png
          Reply  Mon May 7 20:01:14 2018, Rorpheus, Update, General, Use anti-dewhitening + show CARMA/DARMA 
             Reply  Thu May 10 14:13:22 2018, gautam, Update, General, More refinement of DARM control signal projection darmProj.pdfdarmOLTF.pdfDWcomparison.pdf
                Reply  Fri May 11 13:58:42 2018, rana, Update, General, More refinement of DARM control signal projection 
                   Reply  Fri May 11 19:02:52 2018, gautam, Update, General, More refinement of DARM control signal projection darmProj.pdfdarmOLTF.pdfDWcomparison.pdfDARMsuppression.pdfControlLoop.pdf
                      Reply  Sat May 12 10:02:03 2018, rana, Update, General, More refinement of DARM control signal projection 
                         Reply  Sun May 20 17:43:01 2018, rana, Update, Electronics, How to choose resistors 
                      Reply  Mon May 14 18:58:32 2018, Kevin, Update, PonderSqueeze, Squeezing with no SRM homodyne_heatmap.pdfdisplacement_noise.pdfnoise_budget.pdf
                         Reply  Tue Oct 2 23:57:16 2018, gautam, Update, PonderSqueeze, Squeezing scenarios PonderSqueeze_NB_LORIN.pdf
                   Reply  Tue May 15 21:56:57 2018, gautam, Update, General, Stack measurement setup decommissioned 
                      Reply  Thu May 17 09:14:38 2018, Steve, Update, General, Stack measurement setup decommissioned Load_Cell_Measurement_Set_Up.jpgETMX_stack_up_down.png
Message ID: 13846     Entry time: Tue May 15 21:56:57 2018     In reply to: 13833     Reply to this: 13851
Author: gautam 
Type: Update 
Category: General 
Subject: Stack measurement setup decommissioned 

[steve,koji,gautam]

Since we think we already know the stack mass to ~25% (i.e. 5000 +/- 1000 lbs), we decided to restore the ETMX stack. Procedure followed was:

  • Take photos of all dial indicators and spirit level. We were at ~-22 mils on all 3 indicators, with 0 being the level before we touched the stack two Fridays ago, i.e. May4.
  • Raised all four jacks installed underneat blue crossbeams in 5mil increments until we were at +25mils on all of them. At this point, there was negligible load on the load cells on top of the STACIS legs, and we could easily slide the load cells out.
  • Rotated all jack screws clockwise (i.e. moving jack screws downwards) by 270 degrees. The southeast jackscrew was rotated by an additional 360 degrees. This was to undo all the jack-screw raising we did on Friday, May 4.
  • Re-installed jacks which were present originally on the STACIS legs, taking care to center the jack as best as we could by eye on the STACIS leg, per Dennis Coyne's suggestion not to impose shear strain on STACIS legs. There were supposedly never carrying any load, and are according to Steve, are there more for safety purposes.
  • Lowered all four jacks in 5 mil steps until dial indicators read ~0. The Northwest jack resting on the STACIS leg was somehow ~0.5cm (!!) below the blue crossbeam even though the corresponding dial gauge read 0, so we raised the jack until it was barely grazing the bottom of the blue crossbeam (confirmed by looking at the point where the dial indicator started going up again). Not sure why this should have been, best hypothesis we have is that someone (one of us) changed the level of this jack while it was removed from the setup.
  • Checked that jack screws could not be turned by hand. At this point, all the load has to be resting on the jack screws, as the jacks we had installed to raise the blue crossbeams could be slid out from underneath the blue beams and hence were carrying no load.
  • Took photographs of all dial indicators, spirit level. We were satisfied that we had recovered the "nominal" stack alignment as best as we could judge with the available indicators.
  • ETMX Oplev spot had returned to the PD. ETMX watchdog was re-engaged, optic was re-aligned using SLOW bias sliders to center Oplev spot.
  • EX NPRO was turned back on, and the green beam was readily locked to a cavity TEM00 mode yes.

I will upload the photos to the PICASA page and post the link here later.

Quote:
 

In this case, we only need a mass estimate of the end chamber contents with an accuracy of ~25%. If we think we have that already, we don't need to keep doing the jacks-strain gauge adventure.

 

ELOG V3.1.3-